OpenStudy (anonymous):

Find the slope and the y intercept 9-1/4y=10x i know that the slope is 10 (y=mx+b) but I cannot remember how to find the y intercept.

OpenStudy (saifoo.khan):

when y intercpts, x = 0

OpenStudy (anonymous):

yes it has to be an ordered pair so it is (0,??)

OpenStudy (saifoo.khan):

Right.

OpenStudy (saifoo.khan):

insert 0 in the equation instead of x to find that value of y.

OpenStudy (anonymous):

so is it (0,9) ??? I amnot really sure how to do it?

OpenStudy (saifoo.khan):

9-1/4y=10x 9-1/4y=10(0) now solve for y.

OpenStudy (anonymous):

I'm sorry I am lost...dummy terms :D

OpenStudy (saifoo.khan):

y = 1/36

OpenStudy (anonymous):

but how do you get that answer?

OpenStudy (anonymous):

9x4=36? so it would be (0,36) ??

OpenStudy (saifoo.khan):

No. (0, 1/36)

OpenStudy (anonymous):

hmmm. ok thank you for the help I need to practice these

OpenStudy (saifoo.khan):

really.

OpenStudy (saifoo.khan):

uw.